During a rain storm, rain comes straight down with a velocity Vi = -15 m/s and hits the roof of a car perpendicularly. The mass of the rain per second that strikes the roof is 0.060 kg every second. Assuming that the rain comes to rest upon striking the car (Vf = 0 m/s), find the average force exerted by the rain on the roof during a 3 second time interval.
Answer #1

The average force is continuous, so the 3 second time interval is IRRELEVANT. Please explain this to your instructor if necessary!

Average Force = Averate rate of change of momentum = (0.06kg*15m/s)/s = 0.9 N.s/s

Average Force = 0.9 N.s/s = 0.9 N <= ANS

It is WRONG to multiply (or divide!) this result by 3!
Answer #2

F = v*dm/dt

= 15 * 0.06

= 0.9 N

The average force i independen of the time mentioned = 3s